DIDEROT MATHEMATICAL FORUM

A cycle of conferences, called “Diderot Mathematical Forum”, features two conferences a year taking place simultaneously in three European cities exchanging information by telecommunication and addressing in their programmes three different aspects : fundamental mathematics, applications of mathematics and their relation to society (e.g., ethical and epistemological dimensions).

FIRST DIDEROT MATHEMATICAL FORUM

MATHEMATICS AND FINANCE LONDON, MOSCOW, ZURICH

Organisers

London : Dr Mark DAVIS ( Tokyo-Mitsubishi International)

Professor Terry LYONS ( Imperial College, Londres)

Moscou : Professor Albert SHIRYAEV ( Steklov Mathematical Institut, Russian Academy of Sciences)

Professor Alexei ZIZCHENKO ( Department of Mathematics, Russian Academy of Sciences) Professor Boris KASHIN ( Actuarial-Financial Centre, Russian Academy of Sciences)

Zürich : Professor Ole BARNDORFF-NIELSEN ( Aarhus) (for the LATSIS MEETING) Professor Paul EMBRECHTS (ETH)

DMF 1 : MATHEMATICS IN FINANCE LONDON September 24 1996

Organisers Dr Mark DAVIS (TOKYO-MITSUBISHI INTERNATIONAL) Professor Terry LYONS (IMPERIAL COLLEGE)

Location Pippard Lecture Theatre, Level5, The Sherfield Building, Imperial College, South Kensington, London

Programme TUESDAY SEPTEMBER 24 14H -14H05 Professor Adrian Smith (Imperial College)

Welcome 14H05 -14H45 Professor Chris Rogers (Bath)

Just what is the Black and Scholes Formulae? 14H45 -15H25 Dr Riccardo Rebonnato (BZW)

Interest rates, are they infinite dimensional processus? 15H55 -16H35 Dr Sam Howison (Oxford)

Beyond the Black Scholes paradigm; bubbles, crashes and non-linear forecasting non-linéaire 16H35 -17H15 Dr Lane Hughston (Merrill Lynch)

Foreign exchange rates complicate the issues 17H15 -17H55 Professor Michael Dempster (Cambridge)

It is worthless if you cannot compute 18H -19H30 Panel Discussion should enable participants to interact with experts from the financial

industry, regulators, the press and academia.

For more informations:

http://www.ma.ic.ac.uk/diderot.html

DMF 1 : MATHEMATICS AND FINANCE MOSCOW September 24-25 1996

Organisers

Professor Albert SHIRYAEV ( Steklov Mathematical Institut, Russian Academy of Sciences), Professor Alexei ZIZCHENKO ( Department of Mathematics, Russian Academy of Sciences), Professor Boris KASHIN ( Actuarial-Financial Centre, Russian Academy of Sciences)

Location Steklov Mathematical Institut

Programme I. Mathematics, Real Financial World and their Interaction. II. Statistics of the Financial Series and some of their models. III. Some new theoretical results in Financial Mathematics. IV. Pecularities of the Russian Financial Market. V. Educational aspects in Finance and Actuarial Science.

For more informations:

root@shiryaev.msk.ru DMF 1 : FINANCIAL MATHEMATICS, INSURANCE AND SOCIETY LATSIS SYMPOSIUM 1996 ZURICH September 24-25 1996

Organisers Professor Paul EMBRECHTS (ETH) and Dr Rüdiger FREY (ETH)

Location ETH Zentrum, Hauptgebäude, Rämistrasse 101, Room E1.1

Programme TUESDAY SEPTEMBER 24 10H -10H15 Ole Barndorff-Nielsen (International Statistical Institute et Bernoulli Society) and Jean-

Pierre Bourguignon (European Mathematical Society)

Opening address 10H15 -11H15 Hans Bühlmann (ETH)

On value in Insurance and Finance 11H15 -12H15 Hans Föllmer (Berlin)

Interactions between Finance and Probability 14H15 -15H15 Neil Shepard (Oxford)

Interactions between Finance and Statistics 15H15 -16H15 Ragnard Norberg (Copenhague)

After Black-Scholes: What is Insurance? 17H Discussion

WEDNESDAY SEPTEMBER 25 9H -10H David Heath (Cornell)

A characterization of measures of risk 10H -11H Stewart Hodges (Warwick)

New models of security price processes 11H30 -12H30 Richard Olsen (Olsen and Associates)

Views from the Frontier 12H30 Discussion 15H Hans Föllmer (Berlin)

General audience talk ”Strategien gegen den Zufall oder wildes Wetten? Zur Mathematik von Boeersenkursen und Optionen”

16H30 Discussion

For more informations:

Conference Secretary: Mrs G. Baltes tel. +411 632 3400 baltes@math.ethz.ch